OpenStudy (anonymous):

How can i find the x-coordinate of the center of the ellipse with the following equation? (y+4)^2/9+(x-3)^2/18=6

OpenStudy (anonymous):

PLEASE HELP ME SOMEBODY

OpenStudy (anonymous):

(x-h)^2/a^2 + (y-k)^2/b^2 = 1 is the standard form for the equation of an ellipse where (h, k) is the center. based on that info, can you answer your question now?

OpenStudy (anonymous):

the answer was 3 ;; but.. How?

OpenStudy (anonymous):

if we write the equation (y+4)^2/9+(x-3)^2/18=6 in standard form, we get: (x-3)^2/18 + (y+4)^2/9=6 notice h=3... :)